OpenStudy (anonymous):

write the equation 5x-6y-12=0 in the slope intercept form and then find the slope and y intercept of the corresponding line

OpenStudy (anonymous):

slope intercept is Y=MX+B rewriting in slope intercept form we have Y=5/6x-2 slope = 5/6x y intercept =-2

OpenStudy (anonymous):

yes, Using the y=mx+b formula, m=5/6, b=-2 m is the slope and b is the y-intercept